5.) How many milliliters of a 1.50 molar copper (II) sulfate solution contain 25.0 grams of copper (II) sulfate?
Answer #1

molarity = (W/MW)(1000/V in mL)

molarity = 1.50 M

W = 25.0 g

MW = 159.61 g/mol

V = ?

1.50 = (25.0 / 159.61)(1000 / V)

1000/V = 9.58

V = 104.4 mL